# EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering

### v1

@@ 7.3 Performance by Engineering Domain
Figure 7: Domain-level performance comparison of the top four models ranked by mean Reasoning (F1 Score). The plot reveals a “spiky” performance profile, highlighting that model reasoning is highly specialized and not generalized across topics.
7.4 Performance by Problem Difficulty
As expected, model performance degrades as problem complexity increases, but this trend is most pronounced in Final Answer Accuracy. Most models maintain a respectable accuracy of 60-70% on Easy and Intermediate tasks. However, performance drops sharply on Advanced problems, with nearly all models falling into the 40-50% range. For instance, gemini-2.5-pro’s accuracy falls from 68.6% on Intermediate problems to just 45.2% on Advanced.
The Reasoning (F1 Score), by contrast, tells a more complex story, as illustrated in Figure 8. Instead of a uniform decline, the reasoning scores for many top models (such as claude-3-7 and gpt-5) are lowest on Intermediate tasks before recovering slightly on Advanced tasks. A few models (e.g., gemini-2.5-flash and deepseek-reasoner) even show a slight improvement in reasoning as difficulty increases. The most critical finding, however, is that reasoning scores remain universally low (sub-21%) across all three difficulty levels. This suggests the models’ procedural reasoning failures are fundamental and not simply a function of task complexity.
Figure 8: Difficulty-level performance comparison of the top four models ranked by mean Reasoning (F1 Score). The plot shows procedural reasoning performance is critically low across all levels of complexity, not just on advanced tasks.
7.5 Qualitative Error Analysis with LLM-as-Judge
While the quantitative Reasoning (F1 Score) identifies that a reasoning chain is flawed, it cannot explain why. To diagnose the root causes of these failures, we employed our LLM-as-Judge system (described in Section 5) to perform a qualitative error analysis on Gemini 2.5 Pro. We selected this model as our representative subject because it is a strong “all-rounder,” achieving both a top-tier reasoning score (17.97%) and the first rank on all language fluency metrics (BERTScore, ROUGE-2, and ROUGE-L). The analysis was done using GPT-4o (OpenAI, 2024) as the judge or expert evaluator model.
The judge was used to analyze all 4,118 reasoning steps that were flagged by our verifier as a “mismatch” from the ground-truth solution. The results, summarized in Table 2, are striking. The single most significant finding is that the vast majority of flagged steps were not errors at all. Our judge classified 73.94% (3,045 instances) as Alternative Correct or in other words, valid reasoning paths that simply differed from the ground-truth solution. This overwhelmingly indicates that models are frequently penalized by rigid, single-path evaluations for finding a valid, creative, or different solution, a phenomenon masked by low quantitative scores.
For the remaining 1,072 instances identified as genuine errors, Conceptual Error was the dominant failure mode, accounting for 59.1% of all true errors (634 instances). This demonstrates that models are far more likely to fail because they misapply a physical principle or choose the wrong formula, rather than making simple Calculation Errors (14.3%) or Input Errors (7.6%). This finding pinpoints the primary weakness of modern LLMs not in their arithmetic, but in their application of deep, domain-specific knowledge. A full breakdown of error distributions by branch, domain, and difficulty can be found in Appendix M.
Error Category Total Count Percentage (%)
Alternative Correct 3,045 73.94%
Conceptual Error 634 15.40%
Other 204 4.95%
Calculation Error 153 3.72%
Input Error 81 1.97%
Analysis Failed 1 0.02%
Total 4,118 100.0%
Table 2: Overall distribution of error types identified by Gemini 2.5 Pro as LLM-as-Judge across all 4,118 flagged reasoning mismatches.
@@ 8 Conclusion and Future Work
We introduced EngChain, a symbolic benchmark for verifiable, multi-step engineering reasoning. Built from a taxonomy of 90 problem templates across 9 domains and 3 engineering branches, EngChain provides a verifiable evaluation of multi-step engineering problem-solving, assessing the entire reasoning chain rather than only the final answer.
Our comprehensive evaluation of frontier models reveals a stark “right for the wrong reasons” phenomenon: a universal gap between modest Final Answer Accuracy and critically low Reasoning (F1 Score)s. Our qualitative analysis further reveals that this gap is not only due to error, but that a striking majority of flagged mismatches are actually Alternative Correct solutions, which single-path verifiers would wrongly penalize. For genuine failures, we find that Conceptual Errors are the dominant failure mode, not simple calculation mistakes.
Future work will proceed in several key directions. First, we will complete our planned evaluation by benchmarking the science-enhanced, math-enhanced, and general-purpose open models to provide a holistic comparison across all model categories. Second, we will expand the LLM-as-Judge analysis to other top-performing models to create diagnostic profiles of their unique failure modes. Third, to rigorously validate our qualitative methodology, we will conduct two crucial follow-up studies. A Human Evaluation Study will be performed to establish a definitive ground-truth for our judge’s classifications. This will be complemented by a Cross-Judge Consistency study to ensure our error categorizations are not an artifact of our chosen evaluator (GPT-4o (OpenAI, 2024)). This consistency study will involve a diverse set of judges, including a model from a different leading developer, such as Claude 4 Opus (Anthropic, 2025c), and a high-capability open-source model, such as Qwen 2.5 (Qwen, 2024). Finally, upon completion of these studies, all evaluation scripts, prompts, and detailed results will be publicly released to ensure full reproducibility.
@@
First, a significant limitation lies in our primary quantitative metric, the Reasoning (F1 Score). By design, this metric measures alignment against a single, pre-defined ground-truth path. However, our own qualitative analysis (Section 7.5) revealed that a striking majority of steps flagged as “mismatches” were actually Alternative Correct solutions. This indicates that our quantitative F1 score, when used in isolation, significantly underestimates the true reasoning capabilities of the models and inherently penalizes valid, alternative problem-solving approaches. Future work should therefore focus on developing more advanced evaluation methods, such as semantic alignment metrics or dynamic, model-based verifiers, which can programmatically accept and validate multiple correct reasoning paths.
Second, EngChain is entirely synthetic, generated from symbolic templates. While this design ensures verifiability and contamination resistance, it does not capture the linguistic diversity or contextual ambiguity of real-world engineering. More complex, ill-defined problems often require interpreting non-textual information, such as technical diagrams, schematics, or CAD models, which are not part of our current evaluation. This can be addressed in future versions by expanding the benchmark to include multi-modal tasks and by using real-world engineering documents as seeds for semi-structured problem generation.
Third, the benchmark’s scope, while covering 90 problem types, is currently limited to three core engineering branches. It does not yet include other major fields, such as civil, aerospace, or biomedical engineering. A clear next step is to expand the EngChain taxonomy in collaboration with domain experts to cover these additional fields, ensuring a more comprehensive test of engineering knowledge.
@@ 3. Mechanical Engineering
The above prompt was used to query large language models to cross-reference and validate the selection of fundamental engineering domains identified in our literature review.
@@ Appendix B LLM Prompt for Area Validation
The following prompt was used to query large language models to identify the fundamental areas or sub-topics within each core engineering domain.
@@
Appendix C Foundational Textbooks for Problem Selection
The core engineering principles, problem typologies, and realistic constraints for the templates developed by us were sourced from the authoritative textbooks listed in Table 3. These texts, which are standard in top-tier engineering curricula, served as the primary knowledge base to ensure EngChain is pedagogically sound and grounded in the core engineering curriculum. They were also used to validate our selection of fundamental domains and areas.
@@ Appendix D LLM Prompt for Pedagogical Scoring
The following prompt was used to query large language models (specifically, gemini-2.0-flash) to evaluate the pedagogical significance of each sub-topic within its parent domain.
@@
Appendix E Detailed Domain-Aware Parameterization Examples
To ensure that problems are not just mathematically solvable but are also grounded in physical and engineering realism, EngChain constrains all generated values by the principles of the domain being tested. The following provides specific examples of this methodology for each engineering branch.
E.0.1 Chemical Engineering
E.0.2 Electrical Engineering
E.0.3 Mechanical Engineering
Appendix F Authoritative Data Sources for Parameterization
To ensure the physical and engineering realism of EngChain, we built our domain-aware parameterization (discussed in Section 4.4) on authoritative data. We extracted and manually verified extensive lists of physical constants and material properties from the standard handbooks and data sources listed in Table 4.
@@ Appendix F Authoritative Data Sources for Parameterization
Appendix G LLM Prompt for AI-Assisted Quality Assurance (QA)
System Prompt
You are an expert engineering professor and a senior Python developer acting as a peer reviewer for the EngChain benchmark.
@@ 3. Evaluation Rubric & JSON Schema:
Appendix H Multi-Axis Rubric for LLM-Based Template Validation
As part of our AI-Assisted Quality Assurance (QA) workflow, the LLM Peer Reviewer evaluates each template against a set of criteria. The multi-axis rubric used for this validation, detailing each scoring category and its description, is defined in Table 5.
@@ Appendix H Multi-Axis Rubric for LLM-Based Template Validation
Table 5: Multi-axis rubric used for LLM-based template validation. Numerical criteria are rated on a 1–5 scale, and the Human Review Flag is a binary (true/false) indicator.
Appendix I LLM-as-Judge Error Categories
During the second stage of our evaluation framework (Section 5), the LLM-as-a-Judge system performs a qualitative error diagnosis on failed reasoning steps. This provides deeper insight than a quantitative score alone by distinguishing between fundamental reasoning failures and valid alternative solutions. The five distinct error categories used for this classification are defined in Table 6.
Error Category Description
Conceptual Error The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction).
Calculation Error The model used the correct formula but made a mathematical mistake (e.g., $2\times 3=5$).
Input Error The model used the correct formula but pulled the wrong number from the problem context or a previous step.
Alternative Correct The model’s step is valid and logically sound but follows a different method or phrasing.
Other The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors.
Table 6: Error taxonomy used to categorize model mistakes during LLM-as-Judge evaluation.
Appendix J LLM Prompt for Automated Error Analysis (LLM-as-a-Judge)
@@ Appendix I LLM-as-Judge Error Categories
You are an expert engineering professor acting as an automated evaluator. Your task is to analyze the ‘‘MODEL’S STEP" against the ‘‘GROUND-TRUTH STEP" and return a structured JSON object based on the rules and inputs below. CRITICAL RULES: 1. Your ‘‘explanation" MUST logically justify your chosen ‘‘error_category". Do not contradict yourself. For example, do not choose ‘‘Calculation Error" and then state that the calculation is correct. 2. If the model’s step is factually correct but takes a different path than the ground-truth, you MUST use the ‘‘Alternative Correct" category. Do not classify a correct step as ‘‘Other". 3. The final output must be only a raw JSON object. Do not include any introductory text, concluding remarks, or markdown formatting. Error Categories: ‘‘Conceptual Error": The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction). ‘‘Calculation Error": The model used the correct formula but made a mathematical mistake (e.g., 2 * 3 = 5). ‘‘Input Error": The model used the correct formula but pulled the wrong number from the problem context or a previous step. ‘‘Alternative Correct": The model’s step is valid and logically sound, but follows a different method or phrasing than the ground-truth step. ‘‘Other": The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors. Input for Analysis:
@@
Appendix K Evaluated Model Details
This appendix provides a comprehensive overview of all Large Language Models discussed in our experimental framework (Section 6). As outlined in the main text, our evaluation suite is structured around a four-way taxonomy to enable a structured comparison across different training paradigms. These four categories are: (1) Frontier Proprietary LLMs, (2) Science / Engineering Enhanced LLMs, (3) Math Enhanced LLMs, and (4) General Purpose Open LLMs.
Table 7 provides the complete list of models within each category, detailing their developing organization, parameter size (where public), backbone model (for fine-tuned variants), and source identifier (either an API endpoint or a HuggingFace repository path).
@@
Gemini 2.5 Pro Google DeepMind N/A gemini-2.5-pro-2025
Gemini 2.5 Flash Google DeepMind N/A gemini-2.5-flash-2025
DeepSeek V3.2 DeepSeek N/A deepseek-v3.2-20250929
DeepSeek V3.1 DeepSeek N/A deepseek-v3.1-20250922
DeepSeek R1 DeepSeek N/A deepseek-r1-20250120
Grok 4 Heavy xAI N/A grok-4-heavy-0709
Grok 4 Fast xAI N/A grok-4-fast-reasoning
Science / Engineering Enhanced LLMs
Galactica Meta AI 120B facebook/galactica-120b
Minerva Google DeepMind 540B deepmind/minerva
ChemLLM Shanghai AI Lab 13B meta-llama/Llama-2-13B ShanghaiAILab/ChemLLM-13B
SciLit01 TII 7B Qwen/Qwen2-7B-Instruct tiiuae/SciLit-01-7B
SciAgent OpenBioLLM Team 8B mistralai/Mistral-7B-v0.3 OpenBioLLM/SciAgent-8B
@@
WizardMath WizardLM Team 7B mistralai/Mistral-7B-v0.1 WizardLMTeam/WizardMath-7B-V1.1
MetaMath MetaMath Project 7B EleutherAI/llemma-7b meta-math/MetaMath-7B-V1.0
Mathstral Mistral AI 7B mistralai/Mistral-7B-v0.1 mistralai/Mathstral-7B-v0.1
Qwen2.5-Math Qwen Team 7B Qwen/Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
General Purpose Open LLMs
LLaMA 3.1 Meta 8B meta-llama/Llama-3.1-8B
Qwen 2.5 Alibaba Qwen Team 7B Qwen/Qwen2.5-7B-Instruct
Qwen 3 Alibaba Qwen Team 8B Qwen/Qwen3-8B

### v2

@@ 5.3.3 Performance by Engineering Domain
Figure 4: Domain-Level Performance comparison of the top four models ranked by mean Final Answer Accuracy. The plot reveals a “spiky” performance profile, highlighting that model capabilities are highly specialized and not generalized across topics.
5.3.4 Performance by Difficulty Level
Stratifying performance by difficulty (Figure 5) reveals a functional transition in model utility. In the Easy tier, efficiency-oriented architectures like GPT-5 Mini (67.84%) and open models like Llama 3.1 70B (52.55%) perform considerably well, suggesting flagship compute is unnecessary for fundamental engineering problems. However, the Advanced tier exposes a sharp divergence: Llama 3.1 70B collapses to 23.33%, identifying frontier systems as the sole viable option for high-complexity scenarios, led by Gemini 3 Pro (54.55%). Notably, specialized math-enhanced models (e.g., Qwen 2.5 Math) consistently underperform, peaking at just 32.35% on Easy tasks, confirming that abstract mathematical pre-training is insufficient for domain-specific physical reasoning.
Figure 5: Difficulty-Level Performance of the top four models by mean Final Answer Accuracy. The plot shows a “complexity cliff”: efficiency-oriented models degrade sharply at advanced tasks, whereas frontier reasoning models remain stable.
5.4 Error Analysis
To characterize reasoning failures beyond simple accuracy metrics, we analyzed the 17,528 reasoning chains (54.1% of the dataset) that failed Tier 1 verification. This audit reveals that across all models, 85.86% of traces flagged as discrepancies were identified by the AI Tribunal as Alternative Correct derivations, as illustrated in Figure 6. This rate approached saturation in frontier models (97.17%), confirming that advanced models possess a high degree of solution diversity that renders rigid template-matching metrics unsuitable for engineering evaluation. Focusing on confirmed failures, the aggregate data shows that Conceptual Errors (1,318) outnumbered Calculation Errors (711) by a ratio of nearly 2:1. This indicates that while models frequently hallucinate physics principles, they are comparatively more reliable in arithmetic calculations. This trend correlates strongly with complexity: while Easy tasks exhibited a balanced error distribution, Advanced tasks saw the conceptual-to-calculation failure ratio widen to 3.7:1. Finally, 795 “Lucky Guesses” (2.5%), defined as correct answers with structurally invalid reasoning, underscore the necessity of multi-stage verification over final-answer accuracy alone.
Figure 6: Distribution of Identified Reasoning Errors across model categories. While Frontier models primarily utilize valid alternative methods, General Open and Math-Enhanced models exhibit a significantly higher density of confirmed reasoning failures.
@@ 6 Conclusion and Future Work
In this work, we introduced EngTrace, a symbolic benchmark designed to enable verifiable process supervision of physically grounded engineering reasoning in LLMs. Our evaluation of 24 models reveals a “complexity cliff,” where efficiency-oriented and open-weights models degrade steeply in advanced engineering tasks while frontier reasoning models maintain stability. These findings highlight a critical reasoning gap, emphasizing that trustworthy engineering AI requires more than just general logic. Future work will explore tool-augmented reasoning to decouple semantic physical logic from computational precision and investigate uncertainty quantification to determine if models can reliably calibrate confidence in safety-critical engineering scenarios.
@@
First, our current evaluation treats models as isolated inference engines, detached from the practical constraints of applied engineering. By focusing solely on static text generation, we overlook two critical components of reliable design: iterative self-correction and risk assessment. Consequently, while we effectively measure raw theoretical proficiency, our current framework does not fully capture the stability or safety assurance required for deployment in high-stakes engineering workflows.
Second, EngTrace is entirely synthetic, generated from symbolic templates. While this design ensures verifiability and contamination resistance, it does not capture the linguistic diversity or contextual ambiguity of real-world engineering. More complex, ill-defined problems often require interpreting non-textual information, such as technical diagrams, schematics, or CAD models, which are not part of our current evaluation. This can be addressed in future versions by expanding the benchmark to include multi-modal tasks and by using real-world engineering documents as seeds for semi-structured problem generation.
Third, the benchmark’s scope, while covering 90 problem types, is currently limited to three core engineering branches. It does not yet include other major fields, such as civil, aerospace, or biomedical engineering. A clear next step is to expand the EngTrace taxonomy in collaboration with domain experts to cover these additional fields, ensuring a more comprehensive test of engineering knowledge.
Ethical Statement and Broad Impact
This work utilizes synthetic data generated through symbolic templates and standard engineering principles; no private, sensitive, or proprietary industrial data was used. While EngTrace aims to advance reliable engineering reasoning, we emphasize that current LLMs should not be deployed in safety-critical physical systems without rigorous, hardware-in-the-loop verification. We hope this benchmark fosters the development of interpretable, verifiable, and physically grounded AI for high-stakes engineering applications.
Data License
The EngTrace dataset and associated code will be released under the MIT License.
Acknowledgments
We extend our gratitude to the engineering professionals and domain experts for validating the EngTrace framework and reviewing the templates.
@@ Appendix A LLM Prompt for Domain Validation
To identify core domains, we consulted a panel of four frontier models: GPT-4o, Claude 3.5 Sonnet, Gemini 2.5 Pro, and DeepSeek V3. Each was prompted individually to identify cornerstone engineering domains. We finalized the nine core domains via majority voting. The domain identification prompt is provided below.
@@ Appendix B LLM Prompt for Area Validation
Following the domain validation procedure, we utilized the same model panel and aggregated the pedagogical areas via majority voting. The prompt used for area identification is provided below.
@@
Appendix C Foundational Textbooks for Template Selection
The core engineering principles, problem typologies, and realistic constraints for the templates developed by us were sourced from the authoritative textbooks listed in Table 2. These texts, which are standard in top-tier engineering curricula, served as the primary knowledge base to ensure EngTrace is pedagogically sound and grounded in the core engineering curriculum. They were also used to validate our selection of fundamental domains and areas.
@@ Appendix D LLM Prompt for Pedagogical Scoring
The following prompt was used to query LLMs (described in Section A) to evaluate the pedagogical significance of each sub-topic within its parent domain.
@@
Appendix E Authoritative Data Sources for Parameterization
To ensure the physical and engineering realism of EngTrace, we built our domain-aware parameterization on authoritative data. We extracted and manually verified extensive lists of physical constants and material properties from the standard handbooks and data sources listed in Table 3.
@@ Appendix E Authoritative Data Sources for Parameterization
Appendix F Detailed Domain-Aware Parameterization Examples
To ensure that problems are not just mathematically solvable but are also grounded in physical and engineering realism, EngTrace constrains all generated values by the principles of the domain being tested. The following provides specific examples of this methodology for each engineering branch.
F.0.1 Chemical Engineering
F.0.2 Electrical Engineering
F.0.3 Mechanical Engineering
Appendix G LLM Prompt for Automated AI Pre-Screening
You are an expert engineering professor and a senior Python developer acting as a peer reviewer for the EngTrace benchmark.
@@ 3. Evaluation Rubric & JSON Schema:
Appendix H Multi-Axis Rubric for Automated AI Pre-Screening
During automated AI pre-screening, judges in the AI Tribunal evaluate templates against a set of criteria. The evaluation rubric, detailing scoring categories and descriptions, is provided in Table 4.
@@ Appendix H Multi-Axis Rubric for Automated AI Pre-Screening
Table 4: Multi-axis rubric used for AI Tribunal based template validation. Numerical criteria are rated on a 1–5 scale, and the Human Review Flag is a binary (true/false) indicator.
Appendix I Human Certification Interface
To facilitate the rigorous Stage 2 evaluation, we developed a custom web-based annotation tool using the Streamlit framework. This interface, illustrated in Figure 7, was designed to streamline the expert review process while ensuring data consistency for statistical analysis.
Figure 7: The EngTrace Expert Annotation Interface. The dashboard features a dual-view layout for verifying source code against rendered traces, enabling experts to input structured quality scores and definitive admissibility decisions.
The workflow within the interface proceeds as follows:
  1. 1.
    Dual-View Inspection: The main panel displays the rendered natural language Question and step-by-step Solution Trace exactly as they appear in the final dataset. Simultaneously, reviewers can toggle a tab to inspect the underlying Python Source Code, allowing them to verify the algorithmic logic alongside the output.
  2. 2.
    Dynamic Re-sampling: To verify consistency beyond a single static snapshot, the interface allows reviewers to interactively regenerate instances. This enables experts to stress-test the parameterization logic and confirm that physical constraints hold across multiple random seeds.
  3. 3.
    Ordinal Quality Scoring: Reviewers evaluate the template across three dimensions (Physical Plausibility, Mathematical Correctness, and Pedagogical Clarity) using interactive sliders (1–5). These values provide the ordinal data required for calculating Krippendorff’s Alpha ($\alpha$).
  4. 4.
    Binary Certification Gate: The “Certification” toggle forces a definitive Approve or Reject decision, generating the categorical data for Fleiss’ Kappa ($\kappa$).
  5. 5.
    Mandatory Feedback Loop: If a template is rejected, the interface enforces a mandatory text input field for specific feedback (e.g., "The mole balance assumption is invalid for variable volume"). This feedback is logged and routed back to the authors for immediate correction.
Appendix J Template Validation Results
J.1 Stage 1: AI Tribunal Impact
In the initial audit, the AI Tribunal identified 26 Critical Failures and 6 Controversial templates. These templates were manually refactored based on the specific failure modes identified by the models (e.g., unit mismatches or unrealistic material properties) until a 100% pass rate was achieved; representative examples of these identified failures and their associated reasoning are provided in Table 5.
Template Metadata Median Score Votes Model Model Feedback Summary
adiabatic_flame_temperature Chemical Engineering Thermodynamics 2 2/3 GPT-5 Structure clear but temperatures too low; likely $C_{p}$ polynomial or unit mismatch in energy balance.
Claude 4.5 Opus Method sound but values physically inaccurate (e.g., Acetylene $\sim$1493K vs $\sim$2500K); suspects data parameter issues.
Gemini 3 Pro Critical unit error identified in heat capacity integration (incorrect application of gas constant $R$).
multi_segment_rod Mechanical Engineering Mechanics of Materials 3 2/3 GPT-5 Equations correct but ignores yield limits, resulting in unrealistically high stresses.
Claude 4.5 Opus Physics generally sound, but questions the validity of applying large tensile loads to concrete.
Gemini 3 Pro Identifies physically impossible scenarios where brittle materials are subjected to loads exceeding fracture limits.
Table 5: Qualitative Feedback from the AI Tribunal. Examples of critical failures identified during initial screening, categorized by model-specific technical critiques.
J.2 Stage 2: Inter-Annotator Agreement
We measured agreement among domain experts per branch using Fleiss’ Kappa ($\kappa$) for binary admissibility and Krippendorff’s Alpha ($\alpha$) for ordinal quality metrics. As shown in Table 6, we observed perfect agreement ($\kappa=1.0$) across all domains regarding the final inclusion of templates. This unusually high consensus suggests that the AI Tribunal based stage 1 successfully acted as a rigorous filter, ensuring that only high-quality candidates reached the human review stage.
Domain Templates Fleiss’ $\kappa$ Interpretation
Chem. Eng. 30 1.00 Perfect Agreement
Elec. Eng. 30 1.00 Perfect Agreement
Mech. Eng. 30 1.00 Perfect Agreement
Table 6: Binary Admissibility Consensus (Fleiss’ Kappa). Agreement on whether a template should be Included/Rejected.
Krippendorff’s Alpha ($\alpha$)
Domain
Physical
Plausibility
Math.
Correctness
Pedagogical
Clarity
Chem. Eng. 0.224 1.00 0.457
Elec. Eng. 0.372 1.00 0.332
Mech. Eng. 0.647 1.00 0.078
GLOBAL 0.437 1.00 0.269
Table 7: Ordinal Quality Consistency (Krippendorff’s Alpha). Agreement on 1–5 scoring across three axes.
For the ordinal quality ratings, Table 7 reveals a clear distinction between objective and subjective metrics. Mathematical Correctness achieved perfect agreement ($\alpha=1.0$), reflecting the objective nature of verifying derivations. Physical Plausibility showed moderate agreement ($\alpha\approx 0.44$), as experts occasionally debated edge cases in parameter realism. Conversely, Pedagogical Clarity exhibited lower agreement ($\alpha\approx 0.27$), highlighting the inherent subjectivity in evaluating educational "style."
J.3 AI-Human Alignment
We evaluated the performance of the AI Tribunal (Stage 1) against the Human Experts (Stage 2) using two key metrics:
  1. 1.
    Filter Safety (False Positive Rate): The AI Tribunal demonstrated perfect safety (FPR $0.00\%$), as zero templates approved by the AI were subsequently rejected by humans (Target $10\%$ achieved).
  2. 2.
    Scoring Alignment (Spearman’s Rho): The calculated Spearman’s Rho was low ($\rho=0.065,p=0.54$), indicating weak rank correlation. This is likely an artifact of range restriction: since the AI filter was highly effective, the surviving templates all clustered near the maximum quality score (4–5/5). With variance effectively eliminated (a "ceiling effect"), rank correlation becomes statistical noise rather than a meaningful signal of misalignment.
Appendix K Dataset Statistics
Structural Balance and Granularity.
Each branch is represented by exactly 30 templates, ensuring that no single discipline dominates the aggregate performance metrics. This uniform top-level distribution controls for branch-specific bias. At the granular level, the number of Areas (6–7 per branch) reflects the specific curricular density of each discipline; for instance, Mechanical Engineering consolidates multiple sub-topics into broad areas like Fluid Mechanics, resulting in slightly fewer total areas (6) compared to Chemical and Electrical (7), without sacrificing concept coverage.
Difficulty Distribution.
The dataset difficulty is intentionally stratified to mirror curricular progression. We prioritize foundational and applied reasoning (Level 1 & 2: $\approx$75%) to test core engineering laws and procedural workflows, while reserving the most complex synthesis tasks for the advanced tier (Level 3: $\approx$25%). This weighting ensures models are evaluated primarily on standard practice while still distinguishing elite capabilities in multi-physics integration. Table 8 presents the detailed statistical breakdown.
Branch Structure Difficulty Level Total
Dom. Areas Easy Int. Adv.
Chemical 3 7 13 9 8 30
Electrical 3 7 11 12 7 30
Mechanical 3 6 10 13 7 30
Total 9 20 34 34 22 90
Table 8: Dataset Statistics by Branch. The dataset features a perfectly balanced distribution of 30 templates per branch, with difficulty levels scaled to test both foundational knowledge and advanced synthesis.
Appendix L LLM Prompt for Multi-Model Heuristic Verification
@@ Appendix I Human Certification Interface
You are an expert engineering professor acting as an automated evaluator. Your task is to analyze the ‘‘MODEL’S STEP" against the ‘‘GROUND-TRUTH STEP" and return a structured JSON object based on the rules and inputs below. CRITICAL RULES: 1. Your ‘‘explanation" MUST logically justify your chosen ‘‘error_category". Do not contradict yourself. For example, do not choose ‘‘Calculation Error" and then state that the calculation is correct. 2. If the model’s step is factually correct but takes a different path than the ground-truth, you MUST use the ‘‘Alternative Correct" category. Do not classify a correct step as ‘‘Other". 3. The final output must be only a raw JSON object. Do not include any introductory text, concluding remarks, or markdown formatting. Error Categories: ‘‘Conceptual Error": The model applied the wrong scientific principle or formula. ‘‘Calculation Error": The model used the correct formula but made a mathematical mistake. ‘‘Alternative Correct": The model’s step is valid and logically sound, but follows a different method or phrasing than the ground-truth step. ‘‘Other": The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors. Input for Analysis:
@@
Appendix M Evaluation Framework Validation
To ensure that the EngTrace evaluation framework reflects expert engineering judgment, we conducted a rigorous human-alignment study. We established a validation set of 20 instantiated questions, randomly sampled with stratification across all three branches and difficulty levels, and generated responses from five representative models: deepseek-reasoner, gemini-3-pro-preview, gemma-3-27b-it, qwen2.5-math-7b, and Meta-Llama-3.1-8B-Instruct.
This procedure resulted in 100 randomized and anonymized responses. Three subject matter experts per branch independently scored them on a 1–5 Likert scale. Our initial analysis revealed a strong correlation between human-rated reasoning quality and final answer accuracy (Spearman’s $\rho=0.9130$, $p<0.05$), confirming that coherent reasoning traces are highly predictive of numerical correctness. Consequently, we prioritized reasoning process quality as the primary dimension for optimizing our automated evaluator.
M.1 Expert Evaluation Rubrics
The following rubrics were utilized by subject matter experts to evaluate the 100 validation responses. Experts provided independent scores for reasoning and accuracy on a 1–5 Likert scale as detailed in Table 9 and Table 10.
Score Rating Description
5 Excellent Flawless derivation; follows all physical laws and engineering principles.
4 Good Correct reasoning with minor notation inconsistencies or suboptimal paths.
3 Fair Generally sound logic; minor errors in parameter grounding or arithmetic.
2 Poor Significant logical gaps or incorrect governing equations for the domain.
1 Deficient Nonsensical, physically impossible, or complete hallucination of concepts.
Table 9: Expert rubric used for assessing the quality of the reasoning process.
Score Rating Description
5 Excellent Numerically correct within tolerance including correct units.
4 Good Correct value but missing or incorrectly formatted units.
3 Fair Off by a small margin due to rounding; correct magnitude.
2 Poor Numerically incorrect by a significant margin or order of magnitude.
1 Deficient Entirely wrong or absent.
Table 10: Expert rubric for assessing final answer accuracy.
M.2 Ablation Study Results
We systematically assessed various evaluator configurations to optimize alignment with these expert ratings. Table 11 shows the Spearman’s Rank Correlation ($\rho$) between automated pipeline scores and human judgment.
The results demonstrate that the combination of an AI Tribunal ($\tau=0.8$) and numeric step-matching yields the highest alignment ($\rho=0.632$). This alignment significantly degrades upon the removal of the Tribunal ($\rho=0.411$), confirming its necessity for capturing valid alternative derivations. Furthermore, the poor performance of standard NLP metrics ($\rho=0.325$) reinforces the need for the domain-specific numerical verification employed in EngTrace.
Technical Configuration Alignment ($\rho$)
Tribunal ($\tau=0.8$) + Numeric Step-Matching 0.632
Tribunal ($\tau=0.95$) + Numeric Step-Matching 0.581
Tribunal ($\tau=0.6$) + Numeric Step-Matching 0.525
No AI Tribunal; Strict Ground-Truth Matching 0.411
Tribunal + Step-Matching (Numbers ignored) 0.469
ROUGE-L / BERTScore (Standard NLP metrics) 0.325
Table 11: Ablation results showing the Spearman’s rank correlation ($\rho$) between automated reasoning scores and human expert ratings. The data highlights the necessity of the AI Tribunal and numerical verification.
Appendix N Evaluated Model Details
Table 12 details the 24 models evaluated in our experiments (Section 5). We report the developing organization, parameter size, backbone architecture (where applicable), and the specific access identifier used to ensure reproducibility.
@@
Gemini 3 Pro Google N/A gemini-3-pro-001
Gemini 2.5 Pro Google N/A gemini-2.5-pro-002
Gemini 2.5 Flash Google N/A gemini-2.5-flash-002
DeepSeek V3 DeepSeek N/A deepseek-v3
DeepSeek R1 DeepSeek N/A deepseek-reasoner
General Purpose Open LLMs
Llama 3.1 70B Meta 70B meta-llama/Llama-3.1-70B-Instruct
Llama 3.1 8B Meta 8B meta-llama/Llama-3.1-8B-Instruct
Qwen 2.5 72B Alibaba Cloud 72B Qwen/Qwen2.5-72B-Instruct
Qwen 2.5 14B Alibaba Cloud 14B Qwen/Qwen2.5-14B-Instruct
Qwen 2.5 7B Alibaba Cloud 7B Qwen/Qwen2.5-7B-Instruct
Qwen 3 8B Alibaba Cloud 8B Qwen/Qwen3-8B-Instruct
Gemma 3 27B Google 27B google/gemma-3-27b-it
Gemma 2 9B Google 9B google/gemma-2-9b-it
@@
Qwen2.5-Math Qwen Team 7B Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
Mathstral Mistral AI 7B Mistral-7B mistralai/Mathstral-7B-v0.1
WizardMath WizardLM Team 7B Mistral-7B WizardLMTeam/WizardMath-7B-V1.1
MetaMath MetaMath Project 7B Llemma-7B meta-math/MetaMath-7B-V1.0